The market consensus is that Analog Electronic Corporation has an ROE = 15%, a beta of 1.90, and plans to maintain indefinitely its traditional plowback ratio of 1/5. This year’s earnings were $3.20 per share. The annual dividend was just paid. The consensus estimate of the coming year’s market return is 16%, and T-bills currently offer a 5% return.

a. Find the price at which Analog stock should sell.
Answer #1

r = rF + [beta * (Expected Market Return - rF)]

= 5% + [1.9 * (16% - 5%)} = 5% + 20.9% = 25.9%

g = ROE * Plowback Ratio = 15% * (1/5) = 3%

D0 = E0 * (1 - Plowback Ratio) = $3.20 * [1 - (1/5)] = $2.56

P0 = [D0 * (1 + g)] / [r - g]

= [$2.56 * (1 + 0.03)] / [0.259 - 0.03]

= $2.6368 / 0.229 = $11.51